OpenStudy (anonymous):

there are 1644 students in central high school. There are 50 more girls than boys. How many boys are there?

OpenStudy (anonymous):

Divide the total number of students by 2, and then add 25.

OpenStudy (anonymous):

so 1644/2=822 822+25=847 That's the answer.

OpenStudy (anonymous):

wrong smh

OpenStudy (anonymous):

The reason why you would add 25 instead of 50 is because if you add then you'd have to subtract from the girls number. 822-25=797(girls) 822+25=847 Note the 50 difference

OpenStudy (anonymous):

Ooops I flipped the boys and girls excuse me

OpenStudy (anonymous):

797=boys

OpenStudy (welshfella):

let x be number of boys then x + x + 50 = 1644 2x = 1644-50 = 1594 x = 797 girls = 797 + 50 = 847